Solve
4 1/3x + 1 1/2 = 2 1/6

Answer by checkley77(12844)

4 1/3x+1 1/2=2 1/6
13x/3+3/2=13/6
13x/3=13/6-3/2
13x/3=(13-3*3)/6
13x/3=(13-9)/6
13x/3=4/6
13x*6=12
78x=12
x=12/78
x=2/13 answer.